OpenStudy (anonymous):

Molecular Formula Question: Vanillin is the compound containing carbon, hydrogen, and oxygen that gives vanilla beans their distinctive flavor. The combustion of 30.4 mg of vanillin produces 70.4 mg of CO2 and 14.4 mg of H2O. The mass spectrum of vanillin shows a molecular-ion peak at 152 amu. Use this information to determine the molecular formula of vanillin.

OpenStudy (jfraser):

When an organic sample is burned, we get to assume a couple of key things about the combustion: 1) Whatever carbon was in the sample, is now in the \(CO_2\) 2) The MASS of carbon that's now in the \(CO_2\) MUST HAVE COME FROM the sample, and nowhere else 3) the same thing works for the hydrogen in the water. Use the mass percent of carbon in carbon dioxide to find the mass of carbon that came from the sample. For example: \[Mass C = 0.0704gCO_2 * (\frac{12gC}{44gCO_2}) = 0.0192g C\]Do the same thing with the hydrogen. The mass of C and the mass of H will NOT add up to the mass of the sample. The "missing" mass must be the oxygen that was also in the original sample. Take all 3 of those masses, find the moles of each, then divide by the smallest amount to get a whole number mole ratio. This is the empirical formula, which will have its own empirical mass. The molecular formula must have the same ratio of pieces as the empirical formula, but must be some multiple that gets a mass close to the spectrum peak at 152. Find the ratio of\(\frac{MW}{EW}\) and you'll get the multiple of the formula